OpenStudy (anonymous):

What does it mean when a question ask you to describe the nature of the equations roots? this is the equation that came with it : x^2-9x+14=0

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

none

OpenStudy (skullpatrol):

A solution or root of an equation is any value of the variable that turns an open sentence into a true statement. Correct?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

The question ask for you to describe the nature of the equation's roots.

OpenStudy (anonymous):

I know its two but is it two imaginary or real?

OpenStudy (skullpatrol):

What are they, and what is special about them?

OpenStudy (skullpatrol):

$$\Huge x^2-9x+14=0$$ $$\Huge (x-?)(x-?)=0$$

OpenStudy (anonymous):

7,2

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

so their two real?

OpenStudy (skullpatrol):

Yes, they are two positive integers.

OpenStudy (skullpatrol):

$$\Huge{\text{{1, 2, 3, 4,...}}}$$

OpenStudy (anonymous):

i thought negatives could be real answers too?

OpenStudy (skullpatrol):

Yes, they can but not this equation.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (skullpatrol):

$$\Huge (x-7)(x-2)=0$$

OpenStudy (skullpatrol):

" describe the nature of the equations roots "

OpenStudy (skullpatrol):

A solution or root of an equation is any value of the variable that turns an open sentence into a true statement.

OpenStudy (skullpatrol):

Is 0 = 0 a true statement?

OpenStudy (anonymous):

so on 9x^2-24x+16=0 would it be negative or positive four?

OpenStudy (skullpatrol):

$$\Huge 9x^2-24x+16=0 $$ $$\huge (3x)^2-2*12x +4^2=0 $$ $$\huge a^2-2ab+b^2=(a-b)^2$$

OpenStudy (skullpatrol):

$$\Huge (3x-4)^2 = 0$$ Let 3x-4=0 x=4/3